The difference of two positive numbers is 1. The larger number is one less than twice the smaller number. Find the numbers.
Answer by fractalier(6550)

Call the numbers x and y. Then we can write
x - y = 1 and
x = 2y - 1
Now substitute the second into the first and get
2y-1 - y = 1
y - 1 = 1
y = 2
so that
x = 3